# More Modality, More AI: Exploring Design Opportunities of AI-Based Multi-modal Remote Monitoring Technologies for Early Detection of Mental Health Sequelae in Youth Concussion Patients

### v1

@@ 2.3. Designing AI-Powered Systems for Clinical Decision Support
Despite these advancements, the integration of AI-driven decision support systems into clinical workflows remains challenging. One major issue is that these systems often fail to fully consider the needs of stakeholders during the design phase, limiting their practical usability (Cai et al., 2019b; Green and Chen, 2019; Khairat et al., 2018; Lee et al., 2020; Romero-Brufau et al., 2020) and reducing clinicians’ trust (Yang et al., 2016; Elwyn et al., 2013). As a result, AI-driven systems may not effectively assist clinicians in making faster and more accurate decisions (Antoniadi et al., 2021). To address this gap, existing research has sought to incorporate human-centered design principles into AI-powered system development. Studies have shown that involving stakeholders early in the design process and establishing continuous feedback loops (Abdulaal et al., 2021) can enhance the AI system’s integration into clinical workflows (Wang et al., 2023b), which allows AI-powered systems to be tailored to the needs of different stakeholders, ultimately improving the usability of AI system. For example, Yang et al.(Sendak et al., 2020) conducted multiple discussions with clinicians to understand the challenges they face during meetings, collected their feedback, and iteratively refined the design of an AI-powered system based on this input. Similarly, Lee et al. (2020) engaged seven therapists throughout the system design process, which ultimately increased their willingness to use and satisfaction with the system. These studies highlight the importance of human-centered design in AI-driven systems.
Therefore, it is essential to develop a deep understanding of the challenges and needs of clinicians and actively involve them in the design workflow to fully leverage the potential of AI technologies to help clinicians with concussions detect mental health sequelae and make informed clinical decisions.
@@ 3. Formative Study
To better understand the workflow of clinicians treating youth concussion patients, we conducted a formative study focusing on three aspect: (1) how clinicians gather youth patients’ mental health-related information, (2) the decisions that clinicians need to make, and (3) the challenges they face in their current workflow.
@@ 3.1. Study Participants and Procedure
The key expert stakeholders in the targeted scenarios are concussion clinicians with prior experience treating youth concussion patients with mental health symptoms. These highly specialized concussion professionals work in very fast-paced environments and are often overloaded with numerous complex patient cases, which makes their time very limited. For this reason, we recruited six U.S.-based concussion clinicians. Recruitment was facilitated through professional networks within related clinical fields of domain experts in the research team via convenience sampling (Sedgwick, 2013). Six concussion clinicians were recruited as participants for the formative study, where each participant engaged in a semi-structured interview of 35-45 minutes conducted remotely via Zoom. Table 1 provides details on the demographics of the participants and their clinical experience. This study was reviewed and approved by first author’s Institution Institutional Review Board (IRB) and the study complied with the approved procedure for ethical research practices of human subject protection.
At the beginning of each interview, we first asked the participant to recall a recent case where they encountered a youth concussion patient with mental health sequelae. They described how they performed clinical practices to detect mental health sequelae and executed follow-up decision making. While participants were sharing their experiences, we prompted participants to tell more about what type(s) of patient clinical data they collected, used or expected to support the identification of mental health sequelae, both in and outside of clinical settings. After that, participants shared their experience with existing tools, if any, for collecting patient data, monitoring patient status, and detecting mental health sequelae. Participants shared their feedback and insights on the benefits and limitations of existing tools and the expectations of novel tools that could potentially help them. During the interview, we asked participants to refrain from disclosing any personally identifiable information (PII). The complete interview protocol is provided in Appendix A.
All interviews were audio-recorded and transcribed with participants’ consent. We employed the inductive thematic analysis approach (Braun and Clarke, 2012, 2019) to derive key themes. Two researchers first independently coded one session of the interview transcripts to identify meaningful segments and categorize segments into codes. The researchers then discussed individual codes and resolved any discrepancies, developed consensus codes, and applied the resulting codes to the remaining data.
@@ 3.2. Findings
Our qualitative analysis provided a comprehensive understanding of concussion clinicians’ workflow with youth concussion patients, and derived three major challenges that concussion clinicians encountered during the practices: (1) collecting patients’ mental health information outside the clinic; (2) communicating effectively with patients and their family members; (3) Assessing patients’ compliance with clinical recommendations.
3.2.1. Clinical Workflow of Concussion Clinicians
The majority of youth concussion patients are athletes where, and concussions are commonly caused by head injuries during sports participation. After they had a head injury, the youth patients underwent an initial assessment by primary care physicians. According to P4, primary care physicians conducted basic evaluations and symptom scoring based on international guidelines. If specialized care is needed, patients are referred to specialized concussion clinicians.
The clinical tasks of concussion clinicians primarily include seeing new concussion patients and following up with existing patients, which typically follow a similar routine. When patients visit the clinic, concussion clinicians assess the severity of concussion symptoms with standard scales as suggested by clinical guidelines. In addition, concussion clinicians will conduct interview sessions with patients to collect more in-depth information about their health conditions.
”And I guess the only way that I currently assess my patients in that way is by using the symptom score that is published in the international guidelines, at least as of 2017. That’s the symptom score that I use. And then interview. I will very commonly ask about Sleep habits. I will very commonly ask about nutrition habits. I am very aggressive in making I guess my patients go to school.” (P3)
For the majority of concussion patients, the best treatment available is just to keep rest both physically and cognitively. Thus, concussion clinicians will often engage in a conversation with patients after finishing the assessment to provide personalized lifestyle recommendations and self-evaluation methodologies that would be best suited for concussion recovery. For example, one participant (P5) stated that they often provide patients with symptom logs and suggested patients track their daily experiences. Patients will be asked to return the symptom logs during follow-up appointments, which typically occur after one week. Symptom logs are crucial for concussion clinicians because they will heavily rely on the symptom logs to get a thorough understanding of how patients underwent the past week and how the progression of recovery from concussion is. If symptoms persist beyond 28 days, further interventions are initiated, such as referrals to neuropsychologists or specialized concussion clinics (P3, P5, P6).
@@ 3.2.1. Clinical Workflow of Concussion Clinicians
Concussion clinicians (P1, P2, P3, P5, P6) emphasized that mental health concerns are highly prevalent among youth concussion patients. Thus, mental health assessments are an integral part of the concussion assessment process. Concussion clinicians (P6, P4) typically review the patient’s mental health history, including prior diagnoses, current treatments, and medication use. In addition, clinicians leverage standardized questionnaires, such as the Patient Health Questionnaire-9 (PHQ-9) (Levis et al., 2019) and the Generalized Anxiety Disorder 7-item Scale (GAD-7) (Mossman et al., 2017), to identify potential symptoms of mental health issues and the severity of symptoms. Concussion clinicians (P4, P6) also observe signals of emotional distress during the conversation with patients to detect potential mental health issues and will follow up with additional mental health screening if such signals are presented. Clinician (P6) reported that there is a high portion of concussion patients develop anxiety or depression following their injury. In such cases, clinicians (P4) often recommend cognitive behavioral therapy (CBT) interventions as the intervention strategy for the mental health issues mentioned above. However, if mental health symptoms are found to impede concussion recovery, concussion clinicians will need to collaborate with neurologists to determine the intervention plan, such as prescribing appropriate medications. For patients exhibiting severe mental health symptoms, clinicians (P2, P3) will require the patients to complete an additional suicide risk assessment. If no immediate suicide risk is detected, clinicians typically recommend continued rest at home with ongoing symptom monitoring. However, the concussion clinician (P2) mentioned that 10% of youth concussion patients develop suicidal thoughts after a concussion. In such cases, clinicians refer patients to mental health specialists, such as psychiatrists or psychologists.
@@
”In general, of people who I see with like symptoms of like anxiety, depression, things like that after concussion. It’s probably about 40 to 50 percent.” (P6)
3.2.2. Clinicians’ Challenges in Tracking Patients’ Mental Health Information Outside of Clinics
@@
As mentioned above, the first 28 days following a concussion diagnosis is a critical period of time for youth concussion patients’ recovery. However, patients only visit the concussion clinic every one or two weeks, while spending the majority of time outside the clinic, more specifically, outside of clinical supervision. This gap creates significant challenges for concussion clinicians, who (P2, P5) emphasize the importance of monitoring a wide range of mental health symptom indicators during this recovery period. For instance, clinicians (P1, P2, P3, P4, P5) would focus on collecting detailed information about sleep patterns (e.g., sleep hours and wake-up times), and physical activity levels. P3 noted that the sleep data has not really been integrated into concussion statements. Moreover, clinicians (P2, P3, P6) are closely interested in youth patients’ academic performance, such as whether they completed assignments on time, and social behavioral patterns, such as whether they rest at home consistently and maintain social connections rather than withdrawing socially. These physical and cognitive behaviors could provide critical insights for clinicians to assess and reflect on youth patients’ mental health status during concussion recovery.
”…if the worry is getting to the point where they (patients) can’t do their homework… Or if they aren’t leaving the house… not sleeping, really feeling terrible, and it seems to be consuming them… Then I would say that might like cross the line from just typical worry into like, okay, this is more of a mental health anxiety, you know, diagnosis type of thing.” (P2)
@@
Furthermore, P4 stated that during clinical visits, the clinician mainly relies on nonverbal cues in the clinical environment to identify youth patients’ mental health conditions. If any abnormalities are detected, youth patients would be asked directly about mental health-related symptoms, such as suicidal tendencies. Mental health-related information is crucial, as mental health issues impact the clinician’s ability to adjust treatment plans in a timely manner. Moreover, in some cases, youth concussion patients may experience impairment in functioning or panic attacks, and the concussion clinicians will then focus on understanding what they had been through and what events triggered the episode to more accurately evaluate patients’ condition as well as design a more appropriate concussion recovery plan to avoid these negative events happening again, as suggested by P2.
”……I think, having more information about really what was happening at the time, that they had a symptom. What was triggering, that how that we know more real time information could be helpful.” (P2)
@@
Despite the critical need for timely health data of patients at home, concussion clinicians (P2, P5) reported significant challenges in collecting accurate and reliable information about patients’ mental health outside the clinic. Ideally, clinicians would benefit from having real-time access to patients’ health data while they are at home. However, such a system is currently unavailable. Instead, clinicians must rely on patients’ recall of their living experiences during clinical visits to reconstruct what happened during the recovery period. Although these discussions could help inform clinical decision-making, patients’ memory recall is often incomplete or inaccurate, especially when clinical visits span several weeks. This lack of reliable data collection methodology can hinder concussion clinicians’ understanding of the patient’s recovery progress and delay timely mental health interventions, such as referrals or additional assessments. Once youth concussion patients leave the clinic, clinicians have no way to obtain this information in a timely manner.
@@
”…more real-time information could be helpful. Because by the time we see them (patients) in the clinic that could be you know, weeks later they might not even remember that that (mental health sequelae-related symptoms) had happened…… It’s hard for anyone to remember what they did, you know, weeks before.” (P2)
@@
Youth concussion patients currently have some remote communication options, such as making phone calls with the provider team, to report their health status regularly to clinicians when they are outside of clinic. However, clinicians (P2, P3) reported several limitations with phone communication. First, youth concussion patients rarely initiate calls to their concussion clinicians. Moreover, even if youth concussion patients make a phone call, clinicians may not always be in the office or available to answer immediately, which can delay the transmission of critical health information.
@@
An alternative option is using MyChart, a patient portal that allows youth concussion patients to access their medical records and send messages to their healthcare providers. Concussion clinicians (P4) strongly recommend patients use MyChart to share health-related information, particularly when youth concussion patients are reluctant to make direct phone calls. Compared to phone communication, MyChart offers the advantage of asynchronous messaging, enabling clinicians to review patient information and respond at their convenience.
Despite its potential benefits, the adoption of MyChart is limited in current clinical scenarios. Concussion clinicians (P2, P6) noted that youth patients rarely use the messaging feature in MyChart and rarely take the initiative to contact clinicians. Instead, parents of these youth patients are more likely to use MyChart to communicate with clinicians, which impedes the information flow between patients and clinicians.
@@
”One in 10 [patients] maybe, or even one in 20 [patients] might send a message with a question.” (P6)
@@
Furthermore, clinicians caution against relying on MyChart for emergencies, especially when patients experience suicidal tendencies. Clinicians (P2) warned that their busy schedules prevent them from checking messages in real time, which could lead to delays in critical interventions and put patients at risk.
@@
”Yeah, it’s not really immediate. You know, cause they (patients) might send it at night, and then we don’t maybe don’t see it till the next afternoon, and you know. We usually, you know, tell them if it’s an emergency, don’t use that method because we don’t really know when we’re gonna see them or be able to respond to it.” (P2)
@@
3.2.3. Clinicians’ Challenges in Communicating with Patients and Families About Mental Health Issues

### v2

@@ 2.3. Designing AI-based Systems for Clinical Decision Making
Despite these advancements, the integration of AI-based decision support systems into clinical workflows remains challenging. One major issue is that these systems often fail to consider the needs of stakeholders during the design phase, limiting the system’s usability (Cai et al., 2019b; Green and Chen, 2019; Khairat et al., 2018; Lee et al., 2020; Romero-Brufau et al., 2020). Moreover, clinicians remain skeptical about the accuracy of AI-generated predictions, which affects their trust in the system (Bohr and Memarzadeh, 2020; Yang et al., 2016; Elwyn et al., 2013). As a result, AI-based systems may not effectively assist clinicians in making faster and more accurate decisions (Antoniadi et al., 2021).
To address this gap, existing research has sought to incorporate human-centered design principles into AI-based system development. Studies have shown that involving stakeholders early in the design process and establishing continuous feedback loops (Abdulaal et al., 2021) can enhance the integration of AI systems into clinical workflows (Wang et al., 2023b). As a result, AI-based systems can be better aligned with stakeholder needs, ultimately improving their usability. For example, Yang et al.(Sendak et al., 2020) conducted multiple discussions with clinicians to understand the challenges they face during meetings, collected their feedback, and iteratively refined the design of an AI-based system based on this input. Similarly, Lee et al. (Lee et al., 2020) involved seven therapists throughout the design process, which led to greater satisfaction and increased adoption of the system.
These prior works inspired us to begin our project with a formative study to understand concussion clinicians’ challenges and needs. Then, we continued to involve them in the design and development process of the AI-RPM system for the mental health sequelae detection scenario.
@@ 3. Formative Study
To better understand the workflow of clinicians treating youth concussion patients with mental health issues, we conducted a formative study focusing on three main aspects: (1) how clinicians gather youth patients’ mental health-related information, (2) what decisions they need to make, and (3) what the difficulties are in their current workflow.
@@ 3.1. Study Participants and Procedure
The key stakeholders in our scenarios are concussion clinicians with experience treating youth concussion patients with mental health symptoms. These highly specialized concussion professionals work in fast-paced environments and are often overloaded with numerous complex patient cases, which makes their time very limited. For this reason, we were only able to recruit six U.S.-based concussion clinicians who were available to participate in our study. Recruitment was facilitated through professional networks within related clinical fields of domain experts in the research team via convenience sampling (Sedgwick, 2013). Each participant engaged in a semi-structured interview of 35-45 minutes conducted remotely via Zoom. Table 1 provides details on the demographics of the participants and their clinical experience. This study was reviewed and approved by the first author’s institution’s Institutional Review Board (IRB), and the study complied with the approved procedure for ethical research practices of human subject protection.
At the beginning of each interview, we first asked the participant to recall a recent case where they encountered a youth concussion patient with mental health symptoms They described how they performed clinical practices, identified mental health symptoms, and executed follow-up decision-making. While participants were sharing their experiences, we prompted participants to tell more about what type(s) of patient clinical data they collected, used, or expected to support the identification of mental health sequelae, both in and outside of clinical settings. After that, participants shared their experience with existing tools, if any, for collecting patient data, monitoring patient mental health conditions, and identifying mental health sequelae. Moreover, participants shared their insights on the benefits and limitations of existing tools and the expectations of novel tools that could potentially help them. During the interview, we asked participants to refrain from disclosing any personally identifiable information (PII). The complete interview protocol is provided in Appendix A.
All interviews were audio-recorded and transcribed with participants’ consent. We employed the inductive thematic analysis approach (Braun and Clarke, 2012, 2019) to derive key themes. Two researchers first independently coded one session of the interview transcripts to identify meaningful segments and categorize segments into codes. Then, they discussed individual codes, resolved discrepancies, developed consensus codes, and applied the resulting codes to the remaining transcripts. Finally, we iteratively grouped codes into overarching themes and refined these themes through ongoing discussions until we reached a consensus on the final set of themes.
@@ 3.2. Findings
Our qualitative analysis provided a comprehensive understanding of concussion clinicians’ workflow with youth concussion patients and derived three major challenges that concussion clinicians encountered during the practices: (1) tracking patients’ mental health-related information outside the clinic and identifying its severity; (2) communicating effectively with patients and their family members about patients’ mental health issues; (3) assessing patients’ compliance with clinical recommendations.
3.2.1. Integrating Mental Health Screening into Concussion Management
The majority of youth concussion patients are athletes, and concussions are commonly caused by head injuries during sports participation. After they had a head injury, the youth patients underwent an initial assessment by primary care physicians. According to P4, primary care physicians conducted basic evaluations and symptom scoring based on international guidelines. If specialized care is needed, patients are referred to specialized concussion clinicians.
The clinical tasks of concussion clinicians primarily include seeing new concussion patients and following up with existing patients, which typically follow a similar routine. When patients visit the clinic, concussion clinicians assess the severity of concussion symptoms with standard scales as suggested by clinical guidelines. In addition, the clinicians conduct interview sessions with patients to collect more in-depth information about their health conditions.
”I guess the only way that I currently assess my patients in that way is by using the symptom score that is published in the international guidelines…… That’s the symptom score that I use. And then interview. I will very commonly ask about sleep habits. I will very commonly ask about nutrition habits. I am very aggressive in making, I guess my patients go to school.” (P3)
For the majority of concussion patients, the best treatment available is to have a rest both physically and cognitively. Thus, concussion clinicians will have a conversation with patients to provide personalized lifestyle recommendations and self-evaluation methodologies for concussion recovery. For example, P5 stated that they often provide patients with a symptom log and suggested that patients use it to track their daily activities. Patients will be asked to return the symptom logs during follow-up appointments, which typically occur after one week. Symptom logs are essential for concussion clinicians to monitor their patients’ weekly health changes and concussion recovery progress. If symptoms persist beyond 28 days, further interventions are initiated, such as referrals to neuropsychologists or specialized concussion clinics (P3, P5, P6).
@@ 3.2.1. Integrating Mental Health Screening into Concussion Management
Concussion clinicians (P1, P2, P3, P5, P6) emphasized that mental health concerns are highly prevalent among youth patients. Thus, mental health assessments are an integral part of the concussion assessment process. Concussion clinicians (P6, P4) typically review the patient’s mental health history, including prior diagnoses, current treatments, and medication use. In addition, clinicians leverage standardized questionnaires, such as the Patient Health Questionnaire-9 (PHQ-9) (Levis et al., 2019) and the Generalized Anxiety Disorder 7-item Scale (GAD-7) (Mossman et al., 2017), to identify potential symptoms of mental health issues and the severity of symptoms. Concussion clinicians (P4, P6) also observe signals of emotional distress during the conversation with patients to detect potential mental health issues and will follow up with additional mental health screening if such signals are presented. P6 reported that there is a high portion of concussion patients develop anxiety or depression following their injury. In such cases, P4 often recommends cognitive behavioral therapy as the intervention strategy for the mental health issues mentioned above.
However, if mental health symptoms are found to impede concussion recovery, concussion clinicians will need to collaborate with neurologists to determine the intervention plan, such as prescribing appropriate medications. For patients exhibiting severe mental health symptoms, clinicians (P2, P3) will require the patients to complete an additional suicide risk assessment. If no immediate suicide risk is detected, clinicians typically recommend continued rest at home with ongoing symptom monitoring. However, P2 mentioned that 10% of youth concussion patients develop suicidal thoughts after a concussion. In such cases, clinicians refer patients to mental health specialists, such as psychiatrists or psychologists.
@@
“In general, of people who I see with like symptoms of like anxiety, depression, things like that after concussion. It’s probably about 40 to 50 percent.” (P6)
3.2.2. Clinicians’ Struggle to Track and Assess Mental Health Beyond the Clinic
@@
As mentioned above, the first 28 days following a concussion diagnosis is a critical period of time for youth concussion patients’ recovery. However, patients only visit the concussion clinic every one or two weeks, while spending the majority of time outside the clinic, more specifically, outside of clinical supervision. This gap creates significant challenges for concussion clinicians, who (P2, P5) emphasize the importance of monitoring a wide range of mental health symptom indicators during the concussion recovery period. Specifically, clinicians (P1, P2, P3, P4, P5) mentoined the importance of collecting detailed information about sleep patterns (e.g., sleep hours and wake-up times), and physical activity levels. P3 emphasized that the sleep patterns has not really been integrated into clinical systems. Moreover, clinicians (P2, P3, P6) are highly interested in youth patients’ academic performance (whether patients completed assignments on time) and social behavioral patterns (whether patients rest at home consistently and maintain social connections rather than withdrawing socially) during concussion recovery time at home. These academic and social behaviors provide critical insights for clinicians to assess and reflect on the severity of youth patients’ mental health issues during concussion recovery.
“From their (patients) activities, feeling [that they have] little interest or in doing activities that they normally would. Seeing difficulty or decreased performance in either activities or or school.” (P6)
@@
Furthermore, P4 stated that to capture mental health-related anomalies, the clinician mainly relies on nonverbal cues during the clinical visits. If any abnormalities are detected, youth patients would be asked questions directly about their mental health conditions, such as suicidal tendencies, to identify the severity of their mental health symptoms. P5 stated that identifying the severity of the symptoms is crucial for clinicians, as it impacts the clinician’s ability to adjust their treatment plans in a timely manner. Moreover, youth concussion patients may experience impairment in functioning or panic attacks, and concussion clinicians will then focus on understanding what they had been through and what events triggered the episode. In this way clinicians can evaluate patients’ mental health condition more accurately as well as design a more personalized concussion treatment plan, as suggested by P2.
“Anybody who has a high report of anxiety, you know, as one of the symptoms or nervousness that we ask on our concussion symptoms log or if they have an underlying diagnosis of an anxiety, disorder, depression. And then, you know, maybe a little bit more likely to refer them (patients) earlier on.” (P5)
@@
Despite the critical need for timely mental health-related information of patients at home, concussion clinicians (P2, P5) reported significant challenges in accurately and reliably collecting the information outside the clinic. Ideally, clinicians would benefit from having real-time access to patients’ accurate mental health-related information while patients are at home. However, such a system is currently unavailable. Instead, clinicians must rely on patients’ recall of their living experiences during clinical visits to reconstruct what happened during the recovery period. Although these discussions could help inform clinical decision-making, patients’ memory recall is often incomplete or inaccurate, especially when clinical visits span several weeks. This lack of reliable data collection methodology can hinder concussion clinicians’ understanding of the patient’s recovery progress and delay timely mental health interventions, such as referrals or additional assessments. Once youth concussion patients leave the clinic, clinicians have no way to obtain this information in a timely manner.
@@
“…more real-time information could be helpful. Because by the time we see them (patients) in the clinic that could be you know, weeks later they might not even remember that that (mental health sequelae-related symptoms) had happened…… It’s hard for anyone to remember what they did, you know, weeks before.” (P2)
@@
Youth concussion patients currently have some remote communication options. One way is making phone calls with the provider team and reporting patients’ mental health-related information regularly to clinicians when they are outside of the clinic. However, clinicians (P2, P3) reported several limitations with phone communication. First, youth concussion patients rarely initiate calls to their concussion clinicians. Moreover, even if youth concussion patients make a phone call, clinicians may not always be in the office or available to answer immediately, which can delay the transmission of critical health information.
@@
An alternative option is using MyChart, a patient portal that allows patients to access their medical records and send messages to their healthcare providers. Concussion clinicians (P4) strongly recommend that patients use MyChart to share health-related information, particularly when youth concussion patients are reluctant to make direct phone calls. Compared to phone communication, MyChart offers the advantage of asynchronous messaging, enabling clinicians to review patient information and respond at their convenience. Despite its potential benefits, the adoption of MyChart is limited in current clinical scenarios. Concussion clinicians (P2, P6) noted that youth patients rarely use the messaging feature in MyChart and rarely take the initiative to contact clinicians. Instead, parents of these youth patients are more likely to use MyChart to communicate with clinicians.
@@
“One in 10 [patients] maybe, or even one in 20 [patients] might send a message with a question.” (P6)
@@
Furthermore, clinicians caution against relying on MyChart for emergencies, especially when patients experience suicidal tendencies. P2 warned that their busy schedules prevent them from checking messages in real time, which could lead to delays in critical interventions and put patients at risk.
@@
“It’s not really immediate. You know, cause they (patients) might send it at night, and then we don’t maybe don’t see it till the next afternoon, and you know. We usually, you know, tell them if it’s an emergency, don’t use that method because we don’t really know when we’re gonna see them or be able to respond to it.” (P2)
@@
3.2.3. Clinicians’ Struggle to Communicate Mental Health Issues with Patients and Their Families